OpenStudy (anonymous):

Solve \[\sqrt{(x-1)}=\sqrt{(x)}-1\]

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

x-1=x+1?

OpenStudy (anonymous):

No @Hollywood_chrissy

OpenStudy (anonymous):

simply sqaure both sides and then taking the surds part on one side again square it to get the ans ....x-1=x+1-2Vx or 2Vx=2 or Vx=1 or x=1

OpenStudy (pradipgr817):

use (a+b)^2 on the right side and solve it

OpenStudy (anonymous):

sorry I am comfused by the first step.

OpenStudy (pradipgr817):

you square left term it becomes x-1. right.

OpenStudy (anonymous):

yes

OpenStudy (pradipgr817):

to solve the term on right use (a+b)^2. so it becomes ((sqrt(x)-1)^2 = x+1-2sqrt(x)

OpenStudy (anonymous):

huh why?

OpenStudy (pradipgr817):

why not?

OpenStudy (pradipgr817):

so x-1=x+1-2sqrt(x) 2sqrt(x)=2 sqrt(x)=1 x=1

OpenStudy (anonymous):

sorry but I do not understand, why can't we just sqaure the rhs. what is this (a+b)^2?????

OpenStudy (anonymous):

your answer is correct, but I do understand the work.

OpenStudy (anonymous):

*not

OpenStudy (pradipgr817):

when you square the left term it becomes\[(\sqrt{x-1})^{2}=(x-1)^{\frac{ 1 }{ 2} \times 2}=x-1\]

OpenStudy (pradipgr817):

you got this

OpenStudy (anonymous):

yes

OpenStudy (pradipgr817):

now square the right term \[(\sqrt{x}-1)^{2}=(\sqrt{x}-1) \times (\sqrt{x}-1)\] \[x-\sqrt{x} -\sqrt{x}+1\]=\[x-2\sqrt{x}+1\] which can be computed using identity \[(a+b)^{2}=(a+b) \times (a+b) = a^{2}+b^{2}+2ab \]

OpenStudy (pradipgr817):

Is it clear?

OpenStudy (anonymous):

oh I was doing \[\sqrt{x}^2-1^2\]

OpenStudy (pradipgr817):

yes

OpenStudy (anonymous):

thanks @pradipgr817

OpenStudy (pradipgr817):

:) my pleasure

OpenStudy (anonymous):

well go for hit and trial,,1 is a solution CLEARLY !!

OpenStudy (anonymous):

Great